# Towards Flexible and Efficient Diffusion Low Light Enhancer

### v1

@@ Introduction
Reflectance-Aware Trajectory Refinement
Ideally, at every step along the trajectory, the predicted clean image should remain consistent. However, because diffusion models are designed for Gaussian flow, it is challenging for them to maintain consistent prediction results throughout the trajectory, leading to discrepancies during the student’s learning process. Fortunately, given an illumination map $h$ and the noise $z$, the reflectance that shares characteristics with the clean image can be obtained through $x=\frac{y-z}{h}$ based on the Retinex theory. The reflectance information derived from the low-light image is particularly well-suited for incorporating deterministic elements into the teacher model’s trajectory. Building on this insight, we propose the Reflectance-Aware Trajectory Refinement (RATR) module to refine the teacher’s trajectory, thereby reducing fitting errors and providing deterministic guidance to the student model
For illumination estimation, we employ the maximum channel of the low-light image $y$ to represent the estimated illumination map $h^{\prime}$, as the common practice. Regarding ISO noise estimation, similar to the previous non-learning-based denoising method (Buades, Coll, and Morel 2005), the noise can be modeled as the distance between the noisy image and the clean image. We use this distance to estimate the noise map of the input low-light images:
$z^{\prime}=|y-\psi(y)|,$ (7)
where $\psi$ represents a non-learning-based denoising operation, allowing for flexible distillation. With these estimations, we can obtain a latent clean image $\tilde{x}_{0}=\frac{y-z^{\prime}}{h^{\prime}}$. Then, we remap this latent clean image to the diffusion trajectory using the deterministic diffusion process:
$\tilde{x}_{s}=\bar{\alpha}_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}.$ (8)
Then, we refine the trajectory at timestep $s$ with a strength $w_{\text{refine}}$:
$\tilde{x}^{\eta}_{s,u,t}=x^{\eta}_{s,u,t}+w_{refine}(\tilde{x}_{s}-x^{\eta}_{s,u,t}).$ (9)
The refined trajectory is then employed to perform distillation. Extensive experiments in ablation studies show that such refinement greatly improves performance.
Auxiliary Loss
In knowledge distillation for classification, it is widely acknowledged that the student classifier often performs as well as, or even outperforms the teacher classifier. A crucial factor contributing to this success is the direct training signal derived from the data label. However, in LLIE, distillation models tend to exhibit inferior sample quality compared to the teacher. This is primarily because model optimization relies solely on the distillation loss. In ReDDiT, we extend the principles of classification distillation to our model by introducing direct signals from both pixel and feature spaces. In the pixel space, we employ L2 loss on the pixel space:
$\mathcal{L}_{pix}=\lambda_{pix}\|x_{0}-x^{\theta}_{0}(x_{t})\|_{2}^{2}.$ (10)
@@
LIME (Guo, Li, and Ling 2016) (TIP 16) 1 16.760 0.560 0.350 15.240 0.470 0.415 16.880 0.776 0.675
Zero-DCE (Guo et al. 2020) (CVPR 20) 14.861 0.562 0.335 18.059 0.580 0.313 - - -
EnlightenGAN (Jiang et al. 2021) (TIP 21) 17.483 0.652 0.322 18.640 0.677 0.309 16.570 0.734 -
RetinexNet (Wei et al. 2018) (BMVC 18) 16.770 0.462 0.474 18.371 0.723 0.365 17.130 0.798 0.754
DRBN (Yang et al. 2020) (CVPR 20) 19.860 0.834 0.155 20.130 0.830 0.147 23.220 0.927 -
KinD (Zhang, Zhang, and Guo 2019) (ACM MM 19) 20.870 0.799 0.207 17.544 0.669 0.375 16.259 0.591 0.435
KinD++ (Zhang, Zhang, and Guo 2019) (IJCV 20) 21.300 0.823 0.175 19.087 0.817 0.180 - - -
MIRNet (Zamir et al. 2022b) (TPAMI 22) 24.140 0.842 0.131 20.357 0.782 0.317 21.940 0.846 -
LLFlow (Wang et al. 2022a) (AAAI 22) 25.132 0.872 0.117 26.200 0.888 0.137 24.807 0.9193 0.067
SNR-Net (Xu et al. 2022) (CVPR 22) 26.716 0.851 0.152 27.209 0.871 0.157 27.787 0.941 0.054
LLFormer (Wang et al. 2023b) (AAAI 23) 25.758 0.823 0.167 26.197 0.819 0.209 28.006 0.927 0.061
Retinexformer (Cai et al. 2023) (ICCV 23) 27.180 0.850 - 27.710 0.856 - 29.040 0.939 -
PyDiff (Zhou, Yang, and Yang 2023) (IJCAI 23) 4 27.090 0.930 0.100 29.629 0.876 0.194 20.659 0.879 0.275
WCDM (Jiang et al. 2023) (ToG 23) 10 26.947 0.847 0.088 30.461 0.879 0.070 29.852 0.908 0.077
GSAD (Hou et al. 2024) (NIPS 23) 10 27.839 0.877 0.091 28.818 0.895 0.095 28.670 0.944 0.047
@@
SID (Chen et al. 2018) (CVPR 18) 16.97 0.591 23.29 0.703
RetinexNet (Wei et al. 2018) (BMVC 18) 16.48 0.578 20.84 0.617
DeepUPE (Wang et al. 2019) (CVPR 19) 17.01 0.604 21.70 0.662
KinD (Zhang, Zhang, and Guo 2019) (ACM MM 19) 18.02 0.583 21.95 0.672
3DLUT (Zeng et al. 2020) (TPAMI 20) 20.11 0.592 21.66 0.655
DRBN (Yang et al. 2020) (CVPR 20) 19.02 0.577 24.08 0.868
EnGAN (Jiang et al. 2021) (TIP 21) 17.23 0.543 20.02 0.604
Uformer (Wang et al. 2022b) (CVPR 22) 18.54 0.577 23.17 0.859
Restormer (Zamir et al. 2022a) (CVPR 22) 22.27 0.649 25.67 0.827
MIRNet (Zamir et al. 2022b) (TPAMI 22) 20.84 0.605 24.38 0.864
SNR-Net (Xu et al. 2022) (CVPR 22) 22.87 0.625 29.44 0.894
Retinexformer (Cai et al. 2023) (ICCV 23) 24.44 0.680 29.77 0.896
@@
Dataset and Implementation Details
Datasets. We evaluate the performance of our framework across various datasets showcasing noise in low-light image regions, including LOLv1 (Wei et al. 2018), LOLv2 (Yang et al. 2021), SID (Chen et al. 2018), SDSD (Wang et al. 2021), DICM (Lee, Lee, and Kim 2012), LIME (Guo, Li, and Ling 2016), MEF (Li et al. 2018), NPE (Wang et al. 2013), and VV.
SID. The subset of the SID dataset captured by a Sony camera is adopted for evaluation. There are 2697 short-/long-exposure RAW image pairs. The low-/normal-light RGB images are obtained by using the same in-camera signal processing of SID (Chen et al. 2018) to transfer RAW to sRGB. 2099 and 598 image pairs are used for training and testing. SDSD. We adopt the static version of SDSD, captured by a Canon EOS 6D Mark II camera with an ND filter. We use 62:6 and 116:10 low-/normal-light video pairs for training and testing on indoor subsets.
@@
LIME (Guo, Li, and Ling 2016) 5.42 3.84 3.73 4.37 3.63
Zero-DCE (Guo et al. 2020) 3.77 3.61 4.12 6.19 4.81
EnGAN (Jiang et al. 2021) 3.98 4.49 4.95 3.42 3.93
KinD++ (Zhang et al. 2021) 6.19 4.24 3.93 3.44 3.72
SCI (Ma et al. 2022) 3.73 3.52 4.07 6.44 4.05
DCC-Net (Zhang et al. 2022) 3.84 6.27 6.04 7.97 4.66
Retinexformer (Cai et al. 2023) 4.47 3.96 4.12 4.28 4.53
WCDM (Jiang et al. 2023) 6.38 4.62 4.86 4.65 5.39
GSAD (Hou et al. 2024) 4.35 3.97 4.08 5.35 3.18
@@
Table 3: Quantitative comparisons of different methods on DICM, LIME, MEF, NPE, and VV datasets in terms of NIQE, lower values indicate better performance. The best result is highlighted in bold.
@@
Comparison with State-of-the-Art Methods
Results on LOLv1 and LOLv2. On LOLv1 and LOLv2, we compare ReDDiT against LIME (Guo, Li, and Ling 2016), RetinexNet (Wei et al. 2018), KinD (Zhang, Zhang, and Guo 2019), Zero-DCE (Guo et al. 2020), DRBN (Yang et al. 2020), KinD++ (Zhang et al. 2021), EnlightenGAN (Jiang et al. 2021), MIRNet (Zamir et al. 2022b), LLFlow (Wang et al. 2022a), DCC-Net (Zhang et al. 2022), SNR-Net (Xu et al. 2022), LLFormer (Wang et al. 2023b), PyDiff (Zhou, Yang, and Yang 2023), WCDM (Wang et al. 2023a), and GSAD (Hou et al. 2024). Table 1 presents the quantitative results of various LLIE methods, showcasing our method’s superior performance across all compared metrics, including PSNR, SSIM, and LPIPS. New SOTA PSNR results of 28.090, 31.250 and 30.166 can be observed on LoLv1, LoLv2-real and LoLv2-synthetic individually. Notably, ReDDiT consistently outperforms other methods, exhibiting substantial improvements in LPIPS scores, indicative of enhanced perceptual quality. Remarkably, ReDDiT achieves SOTA performance on LOLv2-real/LOLv2-synthetic datasets across all distilled models (8, 4, 2 steps). On LOLv1 datasets, our method attains SOTA results with 8 and 4 steps sampling and remains comparable to previous methods with 2 steps. The visual comparison presented in Figure 2 further highlights the effectiveness of ReDDiT in mitigating artifacts and enhancing image details. Notably, as depicted in the red box of Figure 2, ReDDiT excels in restoring clear structures and intricate details. This underscores our method’s ability to leverage generative modeling to capture natural image distributions and retain such characteristics in the distilled models, resulting in superior visual effects.
Results on SID and SDSD. On SID and SDSD, we compare ReDDiT with SID (Chen et al. 2018), DeepUPE (Wang et al. 2019), Uformer (Wang et al. 2022b), RetinexNet (Wei et al. 2018), KinD (Zhang, Zhang, and Guo 2019), DRBN (Yang et al. 2020), EnlightenGAN (Jiang et al. 2021), MIRNet (Zamir et al. 2022b), SNR-Net (Xu et al. 2022), and Retinexformer (Cai et al. 2023). Table 2 presents the quantitative results on the SID and SDSD, indicating its capability of ReDDiT to handle more complex low-light conditions effectively. Specifically, our method establishes new SOTA PSNR values of 25.32 dB/29.95 dB on SID and SDSD dataset. The visual comparison is provided in supplementary materials.
Results on DICM, LIME, MEF, NPE, and VV. As shown in Figure 3, we directly apply the model trained on the LOLv2-synthetic dataset to these unpaired real-world datasets. From Table 3, it is evident that ReDDiT outperforms all competitors in terms of NIQE scores, demonstrating its robust generalization ability. Clearly, our model adeptly adjusts illumination conditions, effectively enhancing visibility in low-light areas while avoiding over-exposure. This successful balance confirms the superior efficacy of our proposed method.
Comparison with Other Accelerate Methods
@@
The refinement module’s influence becomes more pronounced as the number of sampling steps decreases. In the case of 2-step distillation, Figure 6 illustrates a consistent performance improvement with increasing strength of the refinements. On LOLv2-real, the 4-step distillation results even surpass the performance of the 8-step diffusion model and achieve new SOTA results with $w_{refine}=0.8$. This trend is attributed to the heightened significance of the refinement module in addressing the increasing fitting error of the teacher model as the number of steps decreases. Therefore, the refinement module becomes crucial for accurately estimating the trajectory under such conditions.
@@
Furthermore, the degradation in performance resulting from ablation of either loss term is minor compared to the substantial degradation observed when the Reflectance-Aware Trajectory Refinement Module is removed. This finding highlights the critical role of the Reflectance-Aware Trajectory Refinement Module within our framework.
Efficiency Comparison
We evaluate the efficiency of our method in terms of inference time, frames per second (FPS), and number of parameters, comparing it with recent diffusion-based LLIE methods. The quantitative results of this comparative analysis are presented in Table 4. Our approach, particularly in its 2-step variant, demonstrates superior performance regarding inference speed, FPS, and parameter efficiency. It achieves an optimal balance between computational efficiency and model performance, surpassing other methods in both speed and resource utilization. This comparison highlights the effectiveness of our method in delivering high-performance results with minimal computational overhead
@@
Limitation and Future work
@@
While ReDDiT has excelled in restoration within 2 steps, single-step restoration is not yet optimal, presenting a notable limitation. In single-step restoration, inefficient denoising results in low PSNR and artifacts in detail. Failure cases in the single step will be demonstrated in the supplementary materials. Additionally, the exploration of a lightweight denoising network was not addressed in this work. In our future work, we will continue exploring efficient diffusion-based methods for LLIE. Our central focus will be on investigating the potential of a single-step diffusion model and developing a lightweight denoising network.
Conclusions
@@
In this paper, we introduce ReDDiT, a significant advancement in efficient diffusion models for low-light image enhancement (LLIE). Central to ReDDiT is an innovative distillation technique that integrates deterministic elements and reduces trajectory fitting errors through Reflectance-Aware Trajectory Refinement. These advancements enable ReDDiT to preserve the intrinsic structural integrity of images while reducing sampling steps. ReDDiT achieves performance comparable to previous methods using just 2 steps and sets new state-of-the-art (SOTA) results in 4 and 8 steps. Across 10 benchmark datasets, ReDDiT consistently outperforms existing methods. This represents a promising step toward real-time diffusion models for LLIE, and our research in this area will continue.

### v2

@@ 3.3 Reflectance-Aware Trajectory Refinement
3.4 Auxiliary Loss
In knowledge distillation for classification, direct training signal derived from the data label will help student classifier outperforms the teacher classifier. In ReDDiT, we extend the principles to our model by introducing direct signals from both pixel and feature spaces. In the pixel space, we employ L2 loss on the pixel space:
$\mathcal{L}_{pix}=\lambda_{pix}\|x_{0}-x_{est}\|_{2}^{2}.$ (14)
@@ 4 Experiments
LIME [8] (TIP 16) 1 16.760 0.560 0.350 15.240 0.470 0.415 16.880 0.776 0.675
Zero-DCE [7] (CVPR 20) 14.861 0.562 0.335 18.059 0.580 0.313 - - -
EnlightenGAN [12] (TIP 21) 17.483 0.652 0.322 18.640 0.677 0.309 16.570 0.734 -
RetinexNet [35] (BMVC 18) 16.770 0.462 0.474 18.371 0.723 0.365 17.130 0.798 0.754
DRBN [38] (CVPR 20) 19.860 0.834 0.155 20.130 0.830 0.147 23.220 0.927 -
KinD [45] (ACM MM 19) 20.870 0.799 0.207 17.544 0.669 0.375 16.259 0.591 0.435
KinD++ [45] (IJCV 20) 21.300 0.823 0.175 19.087 0.817 0.180 - - -
MIRNet [43] (TPAMI 22) 24.140 0.842 0.131 20.357 0.782 0.317 21.940 0.846 -
LLFlow [33] (AAAI 22) 25.132 0.872 0.117 26.200 0.888 0.137 24.807 0.9193 0.067
SNR-Net [37] (CVPR 22) 26.716 0.851 0.152 27.209 0.871 0.157 27.787 0.941 0.054
LLFormer [32] (AAAI 23) 25.758 0.823 0.167 26.197 0.819 0.209 28.006 0.927 0.061
Retinexformer [2] (ICCV 23) 27.180 0.850 - 27.710 0.856 - 29.040 0.939 -
PyDiff [48] (IJCAI 23) 4 27.090 0.930 0.100 29.629 0.876 0.194 20.659 0.879 0.275
WCDM [11] (ToG 23) 10 26.947 0.847 0.088 30.461 0.879 0.070 29.852 0.908 0.077
GSAD [10] (NIPS 23) 10 27.839 0.877 0.091 28.818 0.895 0.095 28.670 0.944 0.047
@@
SID [3] (CVPR 18) 16.97 0.591 23.29 0.703
RetinexNet [35] (BMVC 18) 16.48 0.578 20.84 0.617
DeepUPE [28] (CVPR 19) 17.01 0.604 21.70 0.662
KinD [45] (ACM MM 19) 18.02 0.583 21.95 0.672
3DLUT [44] (TPAMI 20) 20.11 0.592 21.66 0.655
DRBN [38] (CVPR 20) 19.02 0.577 24.08 0.868
EnGAN [12] (TIP 21) 17.23 0.543 20.02 0.604
Uformer [34] (CVPR 22) 18.54 0.577 23.17 0.859
Restormer [42] (CVPR 22) 22.27 0.649 25.67 0.827
MIRNet [43] (TPAMI 22) 20.84 0.605 24.38 0.864
SNR-Net [37] (CVPR 22) 22.87 0.625 29.44 0.894
Retinexformer [2] (ICCV 23) 24.44 0.680 29.77 0.896
@@
4.1 Dataset and Implementation Details
Datasets. We evaluate the performance of ReDDiT across various datasets showcasing noise in low-light image regions, including LOLv1 [35], LOLv2 [39], SID [3], SDSD [29], DICM [14], LIME [8], MEF [15], NPE [30], and VV.
@@ 4.1 Dataset and Implementation Details
LIME [8] 5.42 3.84 3.73 4.37 3.63
Zero-DCE [7] 3.77 3.61 4.12 6.19 4.81
EnGAN [12] 3.98 4.49 4.95 3.42 3.93
KinD++ [46] 6.19 4.24 3.93 3.44 3.72
SCI [20] 3.73 3.52 4.07 6.44 4.05
DCC-Net [47] 3.84 6.27 6.04 7.97 4.66
Retinexformer [2] 4.47 3.96 4.12 4.28 4.53
WCDM [11] 6.38 4.62 4.86 4.65 5.39
GSAD [10] 4.35 3.97 4.08 5.35 3.18
@@
Table 3: Quantitative comparisons of different methods on DICM, LIME, MEF, NPE, and VV in terms of NIQE. Lower values indicate better performance, with the best result highlighted in bold.
@@
4.2 Comparison with State-of-the-Art Methods
Results on LOLv1 and LOLv2. On LOLv1 and LOLv2, we compare ReDDiT against LIME [8], RetinexNet [35], KinD [45], Zero-DCE [7], DRBN [38], KinD++ [46], EnlightenGAN [12], MIRNet [43], LLFlow [33], DCC-Net [47], SNR-Net [37], LLFormer [32], PyDiff [48], WCDM [31], and GSAD [10]. Tab. 1 presents the quantitative results of various LLIE methods, showcasing our method’s superior performance across all compared metrics, including PSNR, SSIM, and LPIPS. New SOTA PSNR results of 28.090, 31.250 and 30.166 can be observed on LoLv1, LoLv2-real and LoLv2-synthetic individually. Notably, ReDDiT consistently outperforms other methods, exhibiting substantial improvements in LPIPS scores, indicative of enhanced perceptual quality. Remarkably, ReDDiT achieves SOTA performance on LOLv2-real/LOLv2-synthetic datasets across all distilled models (8, 4, 2 steps). On LOLv1 datasets, our method attains SOTA results with 8 and 4 steps sampling and remains comparable to previous methods with 2 steps. The visual comparison presented in Figure 3 further highlights the effectiveness of ReDDiT in mitigating artifacts and enhancing image details. Notably, as depicted in the red box of Figure 3, ReDDiT excels in restoring clear structures and intricate details. This underscores our method’s ability to leverage generative modeling to capture natural image distributions and retain such characteristics in the distilled models, resulting in superior visual effects.
Results on SID and SDSD. On SID and SDSD, we compare ReDDiT with SID [3], DeepUPE [28], Uformer [34], RetinexNet [35], KinD [45], DRBN [38], EnlightenGAN [12], MIRNet [43], SNR-Net [37], and Retinexformer [2]. Tab. 2 presents the quantitative results on the SID and SDSD, indicating its capability of ReDDiT to handle more complex low-light conditions effectively. Specifically, our method establishes new SOTA PSNR values of 25.32 dB/29.95 dB on SID and SDSD dataset. The visual comparison is provided in supplementary materials.
Results on DICM, LIME, MEF, NPE, and VV. We directly apply the model trained on the LOLv2-synthetic dataset to these unpaired real-world datasets. From Tab. 3, it is evident that ReDDiT outperforms all competitors in terms of NIQE scores, demonstrating its robust generalization ability. As shown in Figure 4, our model adeptly adjusts illumination conditions, effectively enhancing visibility in low-light areas while avoiding over-exposure.
Figure 6: An ablation study conducted to evaluate the choice of $\tilde{x}$. The results indicate that reflectance contributes the most to performance during distillation, while low-light images $y$ offer the second-best performance. Both reflectance and low-light images $y$ effectively explore a better residual space, leading to significant improvements in performance compared to the original Gaussian space ($x_{0}$ as $\tilde{x}$).
4.3 Comparison with Other Accelerate Methods
@@ 4.4 Ablation Study
The refinement module’s influence becomes more pronounced as the number of sampling steps decreases. In the case of 2-step distillation, Figure 8 illustrates a consistent performance improvement with increasing strength of the refinements. On LOLv2-real, the 4-step distillation results even surpass the performance of the 8-step diffusion model and achieve new SOTA results with $\omega=0.8$. This trend is attributed to the heightened significance of the refinement module in addressing the increasing fitting error of the teacher model as the number of steps decreases. Therefore, the refinement module becomes crucial for accurately estimating the trajectory under such conditions.
@@
The ablation on the choice of $\tilde{x}$. We compare the choice of $\tilde{x}$ with refinement using clean images $x_{0}$, low-light images $y$, the enhanced result from the learnable Retinex method (e.g. PairLIE [6]), and our reflectance component, as shown in Figure 6. Both reflectance and low-light images $y$ effectively explore a better residual space, leading to significant improvements in performance compared to the original Gaussian space ($x_{0}$ as $\tilde{x}$). This suggests that shifting the residual space is crucial for enhancing the performance of the student model in LLIE tasks.
Furthermore, the PairLIE results, when used for refinement, perform even poorly compared to refinements using clean images $x_{0}$ and low-light images $y$. Although its prediction is closer to clean images $x_{0}$, it fails to identify a suitable residual space for the student model to learn from, ultimately failing to address the inference gap.
@@
Furthermore, the degradation in performance resulting from ablation of either loss term is minor compared to the substantial degradation observed when the RATR module is removed. This finding highlights the critical role of the RATR module within our framework.
4.5 Efficiency Comparison
We evaluate the efficiency of our method in terms of inference time, frames per second (FPS), and number of parameters, comparing it with recent diffusion-based LLIE methods. The quantitative results of this comparative analysis are presented in Tab. 4. Our approach, particularly in its 2-step variant, demonstrates superior performance regarding inference speed, FPS, and parameter efficiency. It achieves an optimal balance between computational efficiency and model performance, surpassing other methods in both speed and resource utilization. This comparison highlights the effectiveness of our method in delivering high-performance results with minimal computational overhead
@@ 4.5 Efficiency Comparison
4.6 Limitation and Future work
@@
While ReDDiT has excelled in restoration within 2 steps, single-step restoration is not yet optimal, presenting a notable limitation. In single-step restoration, inefficient denoising results in low PSNR and artifacts in detail. Failure cases in the single step will be demonstrated in the supplementary materials. Additionally, the exploration of a lightweight denoising network was not addressed in this work. In our future work, we will continue exploring efficient diffusion-based methods for LLIE. Our focus will be on investigating the potential of a single-step diffusion model and developing a lightweight denoising network.
5 Conclusions
@@
In this paper, we theoretically analyze the two main factors contributing to the performance degradation of the diffusion distillation technique. Building on this, we introduce ReDDiT, a significant advancement in efficient diffusion models for LLIE. Central to ReDDiT is the use of linear exploration in the reflectance-aware residual space, which reduces trajectory fitting errors and the sampling gap. These innovations enable ReDDiT to preserve the intrinsic structural integrity of images while minimizing sampling steps. ReDDiT achieves performance comparable to previous methods with just 2 steps and sets new SOTA results at 4 and 8 steps. Across 10 benchmark datasets, ReDDiT consistently outperforms existing methods. This marks a promising step toward real-time diffusion models for LLIE, and our research in this area will continue.
Supplementary Material
6 Proof of Corollary 1
Given $\tilde{x}_{0}$, it can be represented into the noisy states according to the forward process of diffusion models as follows:
$\tilde{x}_{0}=\frac{x_{t}-\sigma_{t}\tilde{\epsilon}}{\alpha_{t}},$ (17)
Substitute Equation 17 into $\tilde{x}_{s}=\alpha_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}$, we get:
$\tilde{x}_{s}$ $=\alpha_{s}\tilde{x}_{0}+\sigma_{s}\epsilon_{\eta}$ (18)
$=\frac{\alpha_{s}}{\alpha_{t}}x_{t}-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\tilde{\epsilon}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u),$
Given the teacher trajectory $x^{\eta}_{s,u,t}$ defined in the main paper:
$x^{\eta}_{s,u,t}$ $=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$ (19)
$+\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}(\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u))$
$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\epsilon_{\eta}(x_{t},y,t),$
we can obtain the following equation:
$\omega x^{\eta}_{s,u,t}+(1-\omega)\tilde{x}_{s}$ (20)
$=\frac{\alpha_{s}}{\alpha_{t}}x_{t}+\sigma_{s}\epsilon_{\eta}(x_{u},y,u)$
$+\omega\frac{\alpha_{s}}{\alpha_{u}}\sigma_{u}\Big{(}\epsilon_{\eta}(x_{t},y,t)-\epsilon_{\eta}(x_{u},y,u)\Big{)}$
$-\frac{\alpha_{s}}{\alpha_{t}}\sigma_{t}\Big{(}\omega\epsilon_{\eta}(x_{t},y,t)+(1-\omega)\tilde{\epsilon}\Big{)}$
$=\tilde{x}^{\eta}_{s,u,t}.$
7 Implementation details
We train our teacher models for 2 million iterations with the framwork of SR3 on the Gaussian flow [24]. In the distillation stage. and 5000 iterations is required. We employ the Adam optimizer with a learning rate of $1\times 10^{-4}$, applying linear weight decay. To stabilize training, we utilize an exponential moving average (EMA) with a weight of 0.9999 during parameter updates in both the pretrain and distillation stages. We adopt a fine-grained diffusion method with $T=512$ steps and implement a linear noise schedule with endpoints set at $1\times 10^{-4}$ and $2\times 10^{-2}$. The patch size and batch size are set to 96 and 16, respectively. All the experiments are run on a single A100 GPU with 80GB of memory. Training our method with a smaller patch size and batch size on a device with less memory is feasible.
8 Dataset
SID. The subset of the SID dataset captured by a Sony camera is adopted for evaluation. There are 2697 short-/long-exposure RAW image pairs. The low-/normal-light RGB images are obtained by using the same in-camera signal processing of SID [3] to transfer RAW to sRGB. 2099 and 598 image pairs are used for training and testing. SDSD. We adopt the static version of SDSD, captured by a Canon EOS 6D Mark II camera with an ND filter. We use 62:6 and 116:10 low-/normal-light video pairs for training and testing on indoor subsets.
9 More experiment results
Figure 9: Qualitative results on SID and SDSD. Patches highlighted in each image by the red and blue boxes indicate that ReDDiT effectively enhances the visibility, preserves the color, reduces noise, and retains finer details in normal light images. It can be observed that the visual results of our method align more closely with the ground truth. Please zoom in for a clearer view of the image details.
Figure 10: Qualitative results on DICM, LIME, NPE, MEF, and, VV. Our method effectively enhances the visibility and preserves the color. Please zoom in for a clearer view of the image details.
9.1 Visual comparison on SID and SDSD dataset
Due to space constraints in the main manuscript, we only presented quantitative comparisons. Here, we offer additional visual comparisons on the SID [3] and SDSD [29] datasets in Figure 9. These datasets pose significant challenges for low-light image enhancement due to both low light levels and severe noise degradation. Our ReDDiT method demonstrates strong performance in improving brightness and reducing noise on these challenging datasets It can be observed in Figure 9 that our ReDDiT effectively enhances the visibility, preserves the color, improves noise reduction, and retains finer details in normal light images.
9.2 Visual comparison on Unpair dataset
Due to space limitations in the main manuscript, we only provided visual comparisons with GSAD [10]. In Figure 10, we offer additional visual comparisons on DICM, LIME, NPE, MEF, and VV datasets. These results demonstrate the generalizability of ReDDiT to real-world low-light image enhancement scenarios and its consistent performance and superiority in preserving image details and improving visibility, making it a reliable solution for real-world applications.